Draft 14.10, X3J13/93-102. Tue 5-Oct-1993 5:00pm EDT

Programming Language—Common Lisp

A. Appendix

Appendix i ii Programming Language—Common Lisp

A.1 Removed Language Features

A.1.1 Requirements for removed and deprecated features

For this standard, some features from the language described in Common Lisp: The Language have been removed, and others have been deprecated (and will most likely not appear in future Common Lisp standards). Which features were removed and which were deprecated was decided on a case-by-case basis by the X3J13 committee.

Conforming implementations that wish to retain any removed features for compatibility must assure that such compatibility does not interfere with the correct function of conforming programs. For example, symbols corresponding to the names of removed functions may not appear in the the COMMON-LISP package. (Note, however, that this specification has been devised in such a way that there can be a package named LISP which can contain such symbols.)

Conforming implementations must implement all deprecated features. For a list of deprecated features, see Section 1.8 (Deprecated Language Features).

A.1.2 Removed Types

The type string-char was removed.

A.1.3 Removed Operators

The functions int-char, char-bits, char-font, make-char, char-bit, set-char-bit, string-char-p,

and commonp were removed.

The special operator compiler-let was removed.

A.1.4 Removed Argument Conventions

The font argument to digit-char was removed. The bits and font arguments to code-char were removed.

A.1.5 Removed Variables

The variables char-font-limit, char-bits-limit, char-control-bit, char-meta-bit,
char-super-bit, char-hyper-bit, and *break-on-warnings* were removed.

A.1.6 Removed Reader Syntax

The “#,” reader macro in standard syntax was removed.

A.1.7 Packages No Longer Required

The packages LISP, USER, and SYSTEM are no longer required. It is valid for packages with one or more of these names to be provided by a conforming implementation as extensions.

Appendix A–1 A–2 Programming Language—Common Lisp